A gas has a volume of 5.5 liters at 3.3 atm. To expand the volume to 7500 ml, what is the new pressure (in atm) have to be?
Answer #1

We can use the combined gas law to solve this problem, which relates the initial and final conditions of a gas system:

(P1 * V1) / T1 = (P2 * V2) / T2

where P1 and V1 are the initial pressure and volume, respectively, and P2 and V2 are the final pressure and volume, respectively. T1 and T2 represent the initial and final temperatures, but we can assume they are constant for this problem.

Let's plug in the initial conditions first:

P1 = 3.3 atm V1 = 5.5 L

We need to convert the final volume from milliliters to liters:

V2 = 7500 ml = 7.5 L

Now we can solve for P2:

(P1 * V1) / V2 = P2

Plugging in the values we get:

(3.3 atm * 5.5 L) / 7.5 L = 2.42 atm

Therefore, the new pressure has to be 2.42 atm.


answered by: Hydra Master